OpenStudy (anonymous):

find L using the formula ; L = A-2hw over 2h if A = 729.8 in2, h = 17.8 in, and w = 6.4 in ((show work and help solve))

OpenStudy (mandre):

If \[L = \frac{ A-2hw }{ 2h }\] then \[L = \frac{ 729.8-2*17.8*6.4 }{ 2*17.8 }\] That should be pretty simple with a calculator.

OpenStudy (anonymous):

I always end up with an ending equation of 707/35.6 , which i know is wrong but i cant figure out.

OpenStudy (mandre):

2 * 17.8 = 35.6 35.6 * 6.4 = 227.84 729.8 - 227.84 = ???